OpenStudy (anonymous):

make a conjecture about the next term in the sequence. -6,2,-4,-8,-14

OpenStudy (anonymous):

anyone know this answer

OpenStudy (blurbendy):

-4 - 2 = -6 -8 -4 - 2 = -14 -14 -8 -4 - 2 = -28 that would be my guess

OpenStudy (anonymous):

is that supposed to be a -2 or is it really a +2 ??

Directrix (directrix):

I think it is supposed to be a -2 which would possibly yield a -22 for the next term. I'm not sure - waiting to hear if the 2 is a -2.